# Learning tensor networks with parameter dependence for Fourier-based option pricing

### v1

@@ Abstract
A long-standing issue in mathematical finance is the speed-up of pricing options, especially multi-asset options. A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor networks to compress high-dimensional tensors. Another usage of the tensor network is to compress functions, including their parameter dependence. In this study, we propose a pricing method, where, by a tensor learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters. As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices. We show that, in the tested cases involving up to about 10 assets, the proposed method is comparable to or outperforms Monte Carlo simulation with $10^{5}$ paths in terms of computational complexity, keeping the comparable accuracy.
@@ I Introduction
In light of this, applications of quantum-inspired algorithms to option pricing have been also studied, as solutions in the present or the near future [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews. and recently, is utilized also in machine learning [14, 15]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [16, 17, 18, 19] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
In this study, we aim to make FT-based option pricing more efficient using TTs in a more extended way. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We then use these TTs to perform fast option pricing in response to various parameter changes. To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to about ten assets, the computational complexity of our proposed method, which is measured by the number of elementary operations, is either advantageous or comparable to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice. When the computational complexity is comparable, our numerical results indicate that our method is faster in terms of computation time, since our method involves only elementary arithmetic contrary to the Monte Carlo method involving more complicated operations. We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
@@ III.1 Fourier transform-based option pricing
and $r\in\mathbb{R}$ and $\sigma_{1},\cdots,\sigma_{d}>0$ are constant parameters called the risk-free interest rate and the volatilities, respectively. The present time is set to $t=0$ and the present asset prices are denoted by $\vec{S}_{0}=(S_{1,0},\cdots,S_{d,0})$.
@@
where we define $\exp(\vec{x}):=(e^{x_{1}},\cdots,e^{x_{d}})$. $p(\vec{x}|\vec{x}_{0})$ is the probability density function of $\vec{x}:=(\log S_{1}(T),\cdots,\log S_{d}(T))$, the log asset prices at $T$, conditioned on the present value $\vec{x}_{0}=(\log S_{1,0},\cdots,\log S_{d,0})$. In the BS model defined by (5), $p(\vec{x}|\vec{x}_{0})$ is given by the $d$-variate normal distribution:
$p(\vec{x}|\vec{x}_{0})=\frac{1}{\sqrt{(2\pi)^{d}\det\Sigma}}\exp\left(-\frac{1}{2}\left(\vec{x}-\vec{\mu}\right)^{T}\Sigma^{-1}\left(\vec{x}-\vec{\mu}\right)\right),$ (7)
@@ III.3 Monte Carlo-based option pricing
where $\vec{x}_{1},\ldots,\vec{x}_{N_{\rm path}}$ are i.i.d. samples from $p(\vec{x}|\vec{x}_{0})$. On how to sample multivariate normal variables, we leave the detail to textbooks (e.g., [22]) and just mention that it requires more complicated operations than simple multiplications and additions such as evaluations of some elementary functions. Besides, calculating the payoff $v$ with the normal variable $\vec{x}_{i}$ involves exponentiation. In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
IV Learning tensor trains with parameter dependence
@@ IV.1 Outline
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions. The outline is illustrated in Fig. 1. Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in Fig. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$. Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Having these tensor trains, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig. 1 (b). By fixing the index $\vec{k}$ of $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ to the value corresponding to the specified $\vec{p}$ and contracting the adjacent core tensors pairwise, we get the $d$-dimensional TTs $\tilde{\phi}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$. Contracting these TTs as Eq. (15), we get the option price for the specified $\vec{p}$.
@@
Figure 1: Fast option pricing based on TTs proposed in this study. In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD. In (b), we use these TTs to perform fast option pricing for a specified parameter $\vec{p}$.
@@ IV.2 Computational complexity
The computational complexity of TT-based option pricing, involving the contraction of tensors associated with $p_{i}$ and $z_{i}$ and the computation of the inner product between two TTs, is $O(d\chi^{3}N)$ for both computations. In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations. Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec. V.
Here, we ignore the computational complexity of learning the TTs by TCI and SVD and consider that after we get the TTs. This is reasonable if we can use plenty of time to learn the TTs before we need fast option pricing. As discussed in Sec. VI, we can reasonably find such a situation in practice.
@@ V.1 Details
For example, consider the case where $\vec{p}=\vec{\sigma}$. From Eq. (10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{j}z_{j}\lesssim 1$ for each $j$. It is necessary to ensure that the grid points cover the whole of such a region with a sufficiently fine interval. Currently, as $z_{j}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{j}$, $\eta_{j}$ and $\sigma_{j}$ must satisfy $\sigma_{j}\eta_{j}\ll 1$ and $\sigma_{j}\times N\eta_{j}\gg 1$. Therefore, in general, we need the adjustment that $\eta_{j}$ increases when $\sigma$ decreases and vice versa. Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{j}$ set to some constant, and thus we did so. Only in the cases with $d=10$, $\eta_{j}$ is adjusted as
@@
Note that when $\eta_{j}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$. In such a case, the contraction between the core tensor for $z_{j}$ and that for $p_{j}$ is not needed since the latter does not exist, and thus the computational complexity decreases.
@@ V.2 Results
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$. The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively. The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively. Also, the computational times of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$, and subsequently reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{v}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
@@ V.2.1 The case of varying $\vec{\sigma}$
Figure 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. Up to $d=9$, TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ are depicted in Fig. 3(a). By applying SVD to the TTs trained via TCI and subsequently recompressing, the bond dimensions could generally be maintained at 10 or below. The details of this compression by SVD are described in Appendix A. Furthermore, The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
[…]
When the computational complexity of the two approaches is comparable, we argue that TT-based option pricing has an advantage. This is because, in the MC-based method, the unit operations such as random number generation are more complex than those in the TT-based method, which are just multiplications and additions. Indeed, we see that in our demonstration, the TT-based method is in fact advantageous in terms of computational time with the software we use. However, since the computational complexity of the two methods is not significantly different, we should note that the quality of the implementation could affect their computational time.
It has been observed that accuracy tends to decrease as $d$ increases. Consequently, increasing $d$ beyond 10 results in decreased accuracy and makes it difficult to maintain accuracy within that in the MC-based method. At the same time, the bond dimension increases, which makes it difficult to maintain the computational complexity smaller than the MC-based method.
@@
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec]
$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $7.02\times 10^{4}$ $5.0\times 10^{5}$ $3.79\times 10^{-4}$ $2.88\times 10^{-2}$
$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $1.24\times 10^{5}$ $6.0\times 10^{5}$ $5.88\times 10^{-4}$ $3.38\times 10^{-2}$
$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0202 0.0236 $2.08\times 10^{5}$ $7.0\times 10^{5}$ $7.02\times 10^{-4}$ $3.71\times 10^{-2}$
$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0180 0.0253 $2.59\times 10^{5}$ $8.0\times 10^{5}$ $1.46\times 10^{-3}$ $4.09\times 10^{-2}$
$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $3.80\times 10^{5}$ $9.0\times 10^{5}$ $2.51\times 10^{-3}$ $4.51\times 10^{-2}$
$10$ 0.58 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0487 0.0337 $3.72\times 10^{5}$ $1.0\times 10^{6}$ $2.62\times 10^{-3}$ $4.88\times 10^{-2}$
$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0308 0.0357 $1.86\times 10^{6}$ $1.0\times 10^{6}$ $6.08\times 10^{-3}$ $5.11\times 10^{-2}$

### v2

# Learning parameter dependence for Fourier-based option pricing with tensor networks

@@ Abstract
A long-standing issue in mathematical finance is the speed-up of option pricing, especially for multi-asset options. A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor networks to compress high-dimensional tensors. Another usage of the tensor network is to compress functions, including their parameter dependence. Here, we propose a pricing method, where, by a tensor learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters. As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices. We show that, in the tested cases involving up to 11 assets, the proposed method outperforms MC-based option pricing with $10^{5}$ paths in terms of computational complexity while keeping comparable accuracy.
@@ I Introduction
In light of this, applications of quantum-inspired algorithms to option pricing have been also studied, as solutions in the present or the near future [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews. and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17], and differential equations [7, 18]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [19, 20, 17, 21] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
Here, we would like to point out an issue in this TT-based method. That is, we need to rerun TCI to obtain TTs each time the parameters are changed. According to our numerical experiments, TCI takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
@@
To address these issues, we focus on another usage of tensor networks that can embed parameter dependence [22, 23, 24] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We then use these tensor networks to perform fast option pricing in response to various parameter changes. To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice. When the computational complexity is smaller, our numerical results indicate that our method has a further advantage in terms of computation time, since our method involves only elementary arithmetic contrary to the Monte Carlo method involving more complicated operations. We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
@@ III.1 Fourier transform-based option pricing
where $r\in\mathbb{R}$ and $\sigma_{1},\cdots,\sigma_{d}>0$ are constant parameters called the risk-free interest rate and the volatilities, respectively. The present time is set to $t=0$ and the present asset prices are denoted by $\vec{S}_{0}=(S_{1,0},\cdots,S_{d,0})$.
@@
where we define $\exp(\vec{x}):=(e^{x_{1}},\cdots,e^{x_{d}})$. $q(\vec{x}|\vec{x}_{0})$ is the probability density function of $\vec{x}:=(\log S_{1}(T),\cdots,\log S_{d}(T))$, the log asset prices at $T$, conditioned on the present value $\vec{x}_{0}=(\log S_{1,0},\cdots,\log S_{d,0})$. In the BS model defined by (5), $q(\vec{x}|\vec{x}_{0})$ is given by the $d$-variate normal distribution:
$q(\vec{x}|\vec{x}_{0})=\frac{1}{\sqrt{(2\pi)^{d}\det\Sigma}}\exp\left(-\frac{1}{2}\left(\vec{x}-\vec{\mu}\right)^{T}\Sigma^{-1}\left(\vec{x}-\vec{\mu}\right)\right),$ (7)
@@ III.3 Monte Carlo-based option pricing
where $\vec{x}_{1},\ldots,\vec{x}_{N_{\rm path}}$ are i.i.d. samples from $q(\vec{x}|\vec{x}_{0})$. On how to sample multivariate normal variables, we leave the detail to textbooks (e.g., [27]) and just mention that it requires more complicated operations than simple multiplications and additions such as evaluations of some elementary functions. Besides, calculating the payoff $v$ with the normal variable $\vec{x}_{i}$ involves exponentiation. In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
IV Learning parameter dependence with tensor trains
@@ IV.1 Outline
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions. The outline is illustrated in Fig. 1. Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in Fig. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$. We reduce the bond dimensions of these TTs using SVD. Subsequently, we contract adjacent core tensors pairwise and obtain matrix product operators (MPOs). Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Having these MPOs, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig. 1 (b). By fixing the index $\vec{k}$ of $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ to the value corresponding to the specified $\vec{p}$, we get the $d$-dimensional TTs $\tilde{\phi}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$. Contracting these TTs as Eq. (15), we get the option price for the specified $\vec{p}$.
@@
Figure 1: Fast option pricing based on TTs proposed in this study. In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD. Then, we contract the tensors associated with $p_{i}$ and $z_{i}$, resulting in MPOs. In (b), we use these MPOs to perform fast option pricing for a specified parameter $\vec{p}$.
@@ IV.2 Computational complexity
The computational complexity of TT-based option pricing, involving computation of the inner product between two TTs, is $O(d\chi^{3}N)$. In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations. Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec. V.
Here, we ignore the computational complexity of learning the TTs by TCI, SVD, and contraction of tensors associated with $p_{i}$ and $z_{i}$ and consider that after we get the MPOs only. This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing. As discussed in Sec. VI, we can reasonably find such a situation in practice.
@@ V.1 Details
For example, consider the case where $\vec{p}=\vec{\sigma}$. From Eq. (10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{j}z_{j}\lesssim 1$ for each $j$. It is necessary to ensure that the grid points cover the whole of such a region with a sufficiently fine interval. Currently, as $z_{j}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{j}$, $\eta_{j}$ and $\sigma_{j}$ must satisfy $\sigma_{j}\eta_{j}\ll 1$ and $\sigma_{j}\times N\eta_{j}\gg 1$. Therefore, in general, we need the adjustment that $\eta_{j}$ increases when $\sigma$ decreases and vice versa. Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{j}$ set to some constant, and thus we did so. In the cases with $d=10,11$, $\eta_{j}$ is adjusted as
@@
Note that when $\eta_{j}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$.
@@ V.2 Results
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$. The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively. The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively. Also, the computational times of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{v}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively. The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi^{\phi}_{\mathrm{max}}$ and $\chi^{\hat{v}}_{\mathrm{max}}$, respectively, in TABLE 2.
@@ V.2.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in Fig. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{i}$ and $z_{i+1}$ could generally be maintained at less than 10 (refer to TABLE 2). The details of this compression by SVD are described in Appendix A. Furthermore, The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
We argue that, when the computational complexity of the two approaches is at least comparable, the TT-based option pricing has an advantage in terms of computational time. This is because, in the MC-based method, the unit operations such as random number generation are more complex than those in the TT-based method, which are just multiplications and additions. Indeed, in our demonstration, we have shown that the TT-based method is advantageous in terms of computational complexity and exhibits even better performance in terms of computational time. However, it is important to note that the actual wall clock time may vary depending on the quality of the implementation.
It has been observed that accuracy tends to decrease as $d$ increases. Consequently, increasing $d$ further might result in decreased accuracy, making it difficult to maintain accuracy within the MC-based method.
@@
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi^{\phi}_{\mathrm{}}$ $\chi^{v}_{\mathrm{}}$
$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $4.76\times 10^{4}$ $5.0\times 10^{5}$ $3.27\times 10^{-4}$ $2.84\times 10^{-2}$ $4$ $7$
$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $7.94\times 10^{4}$ $6.0\times 10^{5}$ $6.11\times 10^{-4}$ $3.38\times 10^{-2}$ $5$ $7$
$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0202 0.0236 $1.21\times 10^{5}$ $7.0\times 10^{5}$ $7.66\times 10^{-4}$ $3.71\times 10^{-2}$ $6$ $7$
$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0180 0.0253 $1.48\times 10^{5}$ $8.0\times 10^{5}$ $9.07\times 10^{-4}$ $4.09\times 10^{-2}$ $6$ $7$
$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $2.01\times 10^{5}$ $9.0\times 10^{5}$ $1.24\times 10^{-3}$ $4.51\times 10^{-2}$ $7$ $7$
$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0308 0.0357 $2.48\times 10^{5}$ $1.0\times 10^{6}$ $2.19\times 10^{-3}$ $5.11\times 10^{-2}$ $7$ $7$
$11$ auto $3.5\times 10^{-6}$ $3.5\times 10^{-6}$ 0.0341 0.0347 $3.02\times 10^{5}$ $1.1\times 10^{6}$ $2.57\times 10^{-3}$ $5.04\times 10^{-2}$ $8$ $7$